How would you solve the problem 9% of 45
Answer by stanbon(75887)

How would you solve the problem 9% of 45
------
Ans: 0.09*45 = (9/100)45 = (9/20)9 = 81/20 = 4 1/20
================
Cheers,
Stan H.